# Smooth Tests for Normality in ANOVA

### v3

@@ 5 Simulation
In summary, the experiments demonstrate that the proposed tests in Section 3 maintain the nominal significance level well under the null, and exhibit high power under the alternatives, except when $\mathrm{E}[\pi_{k}(Z)]=0$ for all $1\leq k\leq K$. The data-driven tests with the approximated null distribution $H(x)$ reliably control the Type I error and achieve excellent power performance.
Table 11: Empirical rejection rates in Experiment IV, for fixed $K$ ($K=1,\ldots,5$), the data-driven test with limiting $\chi^{2}_{1}$ null distribution ($\hat{K}\&\chi^{2}_{1}$), and the data-driven test with approximated null distribution $H(x)$ ($\hat{K}\&H(x)$).
$H_{0}$ $H_{1}$
$m$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\hat{K}\&\chi^{2}_{1}$ $\hat{K}\&H(x)$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\hat{K}\&\chi^{2}_{1}$ $\hat{K}\&H(x)$
$10$ 0.034 0.044 0.040 0.030 0.028 0.062 0.042 0.040 0.998 0.998 0.996 0.994 0.998 0.998
2020 0.062 0.064 0.060 0.056 0.058 0.060 0.050 0.034 1 1 1 1 1 1
$30$ 0.040 0.040 0.038 0.042 0.048 0.060 0.044 0.034 1 1 1 1 1 1
$40$ 0.038 0.042 0.044 0.028 0.032 0.052 0.040 0.024 1 1 1 1 1 1
$50$ 0.044 0.042 0.042 0.042 0.044 0.060 0.048 0.036 1 1 1 1 1 1
$60$ 0.062 0.080 0.052 0.064 0.074 0.082 0.066 0.016 1 1 1 1 1 1
$70$ 0.062 0.066 0.060 0.050 0.062 0.052 0.042 0.050 1 1 1 1 1 1
$80$ 0.046 0.048 0.042 0.058 0.036 0.040 0.038 0.042 1 1 1 1 1 1
$90$ 0.060 0.058 0.050 0.052 0.046 0.052 0.042 0.032 1 1 1 1 1 1
$100$ 0.052 0.038 0.046 0.058 0.042 0.054 0.048 0.018 1 1 1 1 1 1
$110$ 0.040 0.034 0.038 0.032 0.040 0.052 0.046 0.032 1 1 1 1 1 1
$120$ 0.056 0.044 0.066 0.046 0.046 0.066 0.058 0.040 1 1 1 1 1 1
$130$ 0.052 0.058 0.054 0.056 0.050 0.058 0.052 0.028 1 1 1 1 1 1
$140$ 0.050 0.060 0.054 0.042 0.040 0.074 0.068 0.028 1 1 1 1 1 1
$150$ 0.056 0.064 0.054 0.058 0.060 0.054 0.052 0.028 1 1 1 1 1 1
Table 12: Empirical frequency of $\hat{K}$ under $H_{0}$ in Experiment IV
$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
$10$ 0.972 0.024 0.004 0 0
2020 0.986 0.012 0 0.002 0
$30$ 0.980 0.018 0.002 0 0
$40$ 0.992 0.008 0 0 0
$50$ 0.990 0.010 0 0 0
$60$ 0.996 0.004 0 0 0
$70$ 0.994 0.006 0 0 0
$80$ 0.996 0.004 0 0 0
$90$ 0.996 0.004 0 0 0
$100$ 0.990 0.010 0 0 0
$110$ 0.994 0.006 0 0 0
$120$ 0.992 0.006 0.002 0 0
$130$ 0.992 0.008 0 0 0
$140$ 0.986 0.014 0 0 0
$150$ 0.998 0.002 0 0 0
Table 13: Empirical frequency of $\hat{K}$ under $H_{1}$ in Experiment IV
$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
$10$ 0 0.990 0.008 0.002 0
2020 0 0.982 0.018 0 0
$30$ 0 0.990 0.008 0 0.002
$40$ 0 0.994 0.006 0 0
$50$ 0 0.996 0.004 0 0
$60$ 0 0.996 0.004 0 0
$70$ 0 0.994 0.006 0 0
$80$ 0 0.996 0.004 0 0
$90$ 0 1 0 0 0
$100$ 0 0.998 0.002 0 0
$110$ 0 0.996 0.004 0 0
$120$ 0 1 0 0 0
$130$ 0 1 0 0 0
$140$ 0 0.988 0.012 0 0
$150$ 0 1 0 0 0
6 Discussion of other ANOVA models
In this section, we briefly discuss smooth tests for normality of other ANOVA models beyond those in Section 3. The testing approach above is also applicable to these models.
6.1 One-way fixed effects model with heterogeneous means and variances
We first consider the one-way ANOVA model with potentially different group means and variances, given by
$Y_{ij}=\mu_{j}+\varepsilon_{ij}=\mu_{j}+\sigma_{j}e_{ij},\quad i=1,\ldots,N_{j},\quad j=1,\ldots,J.$ (6.1)
Here $\{\mu_{j}\}_{j=1}^{J}$ and $\{\sigma_{j}\}_{j=1}^{J}$ denote the group-specific means and standard deviations. Owing to heterogeneity across groups, the estimators of each group mean and variance rely solely on the corresponding observations, in the sense that $\hat{\mu}_{j}=N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij},\hat{\sigma}_{j}^{2}=N_{j}^{-1}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\hat{\mu}_{j}\right)^{2}$ for $j=1,\ldots,J$. Accordingly, we define $\hat{\varepsilon}_{ij}=Y_{ij}-\hat{\mu}_{j}$, $\hat{e}_{ij}=\hat{\varepsilon}_{ij}/\hat{\sigma}_{j}=(Y_{ij}-\hat{\mu}_{j})/\hat{\sigma}_{j}$, and $\hat{Z}_{ij}=\Phi(\hat{e}_{ij})$. To implement the smooth test, one may follow the procedure in Section 3 and derive an expression for $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$, $k=1,\ldots,K$, that accounts for the estimation effects. The test statistic is then constructed as a quadratic form of $N^{-1/2}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}(\hat{Z}_{ij})$, normalized by the inverse of its asymptotic covariance matrix. Under mild conditions, an asymptotic $\chi^{2}$ test can be established for model (6.1), thereby extending the unified framework to this case.
6.2 One-way random effects model
Next, we consider the one-way random effects model:
$Y_{ij}=\mu+A_{j}+\varepsilon_{ij}=\mu+A_{j}+\sigma e_{ij},\quad i=1,\ldots,N_{j},\quad j=1,\ldots,J,$ (6.2)
where the random effects $\{A_{j}\}_{j=1}^{J}$ are i.i.d. with $\mathrm{E}[A_{j}]=0$, $\operatorname{Var}[A_{j}]=\sigma_{a}^{2}$, and are independent of the errors $\{e_{ij}\}_{i=1,j=1}^{N_{j},J}$. We estimate $\sigma^{2}$ by
$\hat{\sigma}^{2}=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\bar{Y}_{\cdot j}\right)^{2}=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\hat{\varepsilon}_{ij}^{2},$
where $\bar{Y}_{\cdot j}=N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij}$ and $\hat{\varepsilon}_{ij}=Y_{ij}-\bar{Y}_{\cdot j}$. Note that $\bar{Y}_{\cdot j}$ coincides with $\hat{\mu}_{j}$ in Section 3.2, so the expression for $\hat{\sigma}^{2}$ is identical to that in Section 3.2. Consequently, the methodology in Section 3.2 can be directly applied to test the normality of $e_{ij}$ in (6.2), since the random effects $A_{j}$ are removed by within-group centering.
6.3 Two-way ANOVA models
We now extend the discussion to two-way ANOVA models. Consider the two-way fixed effects model:
$Y_{ij_{1}j_{2}}=\mu_{j_{1}j_{2}}+\varepsilon_{ij_{1}j_{2}}=\big(\mu+\alpha_{j_{1}}+\beta_{j_{2}}+\gamma_{j_{1}j_{2}}\big)+\sigma e_{ij_{1}j_{2}},$ (6.3)
subject to the identifiability constraints
$\sum_{j_{1}=1}^{J_{1}}\alpha_{j_{1}}=\sum_{j_{2}=1}^{J_{2}}\beta_{j_{2}}=\sum_{j_{1}=1}^{J_{1}}\gamma_{j_{1}j_{2}}=\sum_{j_{2}=1}^{J_{2}}\gamma_{j_{1}j_{2}}=0.$
Inspired by the methodology developed in Section 3, we can construct analogous estimators for $\mu$, $\{\alpha_{j_{1}}\}_{j_{1}=1}^{J_{1}}$, $\{\beta_{j_{2}}\}_{j_{2}=2}^{J_{2}}$, $\{\gamma_{j_{1}j_{2}}\}_{j_{1}=1,j_{2}=1}^{J_{1},J_{2}}$, and $\sigma^{2}$, and derive the corresponding test statistic. The theoretical framework in Section 3 remains valid in this setting.
Similarly, consider the two-way random effects model:
$Y_{ij_{1}j_{2}}=\mu_{j_{1}j_{2}}+\varepsilon_{ij_{1}j_{2}}=\big(\mu+A_{j_{1}}+B_{j_{2}}+D_{j_{1}j_{2}}\big)+\sigma e_{ij_{1}j_{2}},$ (6.4)
where the i.i.d. random effects $\{A_{j_{1}}\}$, $\{B_{j_{2}}\}$, $\{D_{j_{1}j_{2}}\}$ and the errors $\{e_{ij}\}$ are mutually independent. To test the normality of $e_{ij}$ in (6.4), one can apply the strategy of Section 6.2, first eliminating the random effects by within-cell centering and then constructing the test statistic based on the residuals.
To conclude, the methodology proposed in Section 3 can be seamlessly extended to more general ANOVA settings. This unified framework offers a flexible and effective approach to testing for normality across diverse linear models.
@@ 7 Concluding remarks
In this paper, we establish Neyman’s smooth tests for assessing the normality assumption in ANOVA models. For three types of one-way fixed effects models, we derive the asymptotic properties of the proposed tests and validate their finite-sample performance through extensive simulations. Our results provide a rigorous and practical tool for evaluating normality in a broad range of ANOVA settings.
Several directions remain open for future research. First, while the proposed tests are developed for specific ANOVA models, the true data-generating mechanism may not be known in practice. A natural extension is to combine our procedures with preliminary structural tests to identify the most appropriate ANOVA specification before applying the proposed methodology. Second, the assumption of independent random errors can be relaxed. For example, repeated clinical trials often involve correlated errors (Ma et al.,, 2012; Schober and Vetter,, 2018; Langenberg et al.,, 2022). In such cases, our smooth test framework may be adapted by incorporating estimated correlation structures. Finally, for the data-driven selection rule, it is of theoretical interest to explore the scenario where the upper bound $D=D(N)$ diverges with $N$, that is, $D\to\infty$ as $N\to\infty$.
Appendix A Appendix
We provide the proof of our main theoretical results, along with some additional numerical results, in this Appendix. The symbol $\lesssim$” means that the left side is bounded by a positive constant times the right side. The symbol $\asymp$” means that both sides are asymptotically equivalent.
Proof of Theorem 1.
We first study the estimation effects of $\hat{\mu}-\mu$ and $\hat{\sigma}^{2}-\sigma^{2}$. Note that
$\hat{\mu}-\mu=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\mu\right)=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\varepsilon_{ij}=\frac{\sigma}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}.$ (A.1)
Under $H_{0}$, $e_{ij}\sim\mathcal{N}(0,1)i.i.d$., then $\sqrt{N}(\hat{\mu}-\mu)\sim\mathcal{N}(0,\sigma^{2})$, and $\hat{\mu}-\mu=O_{p}(N^{-1/2})$ as $N\to\infty$. Besides,
$\hat{\sigma}^{2}-\sigma^{2}=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\left(Y_{ij}-\hat{\mu}\right)^{2}-\sigma^{2}\right)$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\varepsilon_{ij}^{2}-\sigma^{2}\right)-\left(\hat{\mu}-\mu\right)^{2}$

### v4

@@ 5 Simulations
To further assess robustness with respect to the number of groups $J$, we also consider a variant of Experiment III with $J=10$ (denoted as Experiment III). The results corroborate our main conclusions; see Section A.10 of the Online Appendix for details. In summary, the experiments demonstrate that the proposed smooth tests in Section 3 maintain the nominal significance level well under the null, and exhibit high power under the alternatives except when $\mathbb{E}[\pi_{k}(Z)]=0$ for all $1\leq k\leq K$. The data-driven tests with the approximated null distribution $H(x)$ reliably control the Type I error and achieve excellent power performance. In contrast with the three classical methods, our approach is supported by solid theoretical guarantees and demonstrates uniformly stable and competitive performance across all settings considered.
Table 7: Empirical rejection rates under $H_{0}$ in Experiment III. The empirical power under $H_{1}$ equals 1 for all sample sizes and all considered tests.
Smooth tests SW JB KS
$m$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\widehat{K}\&\chi^{2}_{1}$ $\widehat{K}\&H(x)$
10 0.050 0.046 0.048 0.048 0.052 0.076 0.050 0.046 0.038 0
20 0.072 0.060 0.050 0.046 0.050 0.054 0.048 0.064 0.048 0
30 0.054 0.046 0.054 0.038 0.056 0.074 0.062 0.064 0.062 0
40 0.046 0.048 0.060 0.054 0.048 0.054 0.050 0.050 0.046 0
50 0.044 0.060 0.052 0.052 0.048 0.040 0.038 0.042 0.050 0
60 0.066 0.044 0.048 0.052 0.058 0.048 0.042 0.048 0.052 0
70 0.044 0.048 0.038 0.046 0.048 0.072 0.064 0.052 0.044 0
80 0.050 0.044 0.048 0.038 0.042 0.074 0.068 0.042 0.044 0
90 0.040 0.060 0.046 0.054 0.046 0.048 0.042 0.058 0.046 0
100 0.062 0.056 0.042 0.050 0.048 0.058 0.052 0.048 0.044 0
110 0.060 0.046 0.066 0.052 0.050 0.068 0.054 0.058 0.060 0
120 0.034 0.044 0.038 0.038 0.036 0.058 0.056 0.046 0.052 0
130 0.038 0.050 0.042 0.052 0.050 0.058 0.048 0.052 0.056 0
140 0.038 0.050 0.048 0.042 0.044 0.044 0.040 0.058 0.068 0
150 0.064 0.056 0.048 0.058 0.058 0.046 0.042 0.046 0.048 0
Table 8: Empirical frequency of $\widehat{K}$ in Experiment III
Under $H_{0}$ Under $H_{1}$
$m$ $\widehat{K}=1$ $\widehat{K}=2$ $\widehat{K}=3$ $\widehat{K}=4$ $\widehat{K}=5$ $\widehat{K}=1$ $\widehat{K}=2$ $\widehat{K}=3$ $\widehat{K}=4$ $\widehat{K}=5$
10 0.976 0.024 0 0 0 0.032 0.004 0 0.418 0.546
20 0.974 0.018 0.006 0 0.002 0 0 0 0.314 0.686
30 0.986 0.014 0 0 0 0 0 0 0.174 0.826
40 0.990 0.010 0 0 0 0 0 0 0.098 0.902
50 0.998 0.002 0 0 0 0 0 0 0.064 0.936
60 0.988 0.012 0 0 0 0 0 0 0.036 0.964
70 0.992 0.008 0 0 0 0 0 0 0.022 0.978
80 0.990 0.008 0.002 0 0 0 0 0 0.018 0.982
90 0.988 0.012 0 0 0 0 0 0 0.008 0.992
100 0.996 0.004 0 0 0 0 0 0 0.008 0.992
110 0.994 0.004 0.002 0 0 0 0 0 0.002 0.998
120 0.992 0.006 0.002 0 0 0 0 0 0 1
130 0.998 0.002 0 0 0 0 0 0 0 1
140 0.992 0.008 0 0 0 0 0 0 0.002 0.998
150 0.996 0.004 0 0 0 0 0 0 0 1
Figure 3: The sample means and error bars of $\widehat{K}$ in Experiment III
Figure 4: The sample means and error bars of $\widehat{K}$ in Experiment IV
Table 9: Empirical rejection rates under $H_{0}$ in Experiment IV
Smooth tests SW JB KS
$m$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\widehat{K}\&\chi^{2}_{1}$ $\widehat{K}\&H(x)$
10 0.034 0.044 0.040 0.030 0.028 0.062 0.042 0.046 0.008 0
20 0.062 0.064 0.060 0.056 0.058 0.060 0.050 0.038 0.026 0.002
30 0.040 0.040 0.038 0.042 0.048 0.060 0.044 0.042 0.032 0
40 0.038 0.042 0.044 0.028 0.032 0.052 0.040 0.042 0.034 0
50 0.044 0.042 0.042 0.042 0.044 0.060 0.048 0.052 0.048 0
60 0.062 0.080 0.052 0.064 0.074 0.082 0.066 0.056 0.044 0
70 0.062 0.066 0.060 0.050 0.062 0.052 0.042 0.040 0.046 0
80 0.046 0.048 0.042 0.058 0.036 0.040 0.038 0.044 0.046 0
90 0.060 0.058 0.050 0.052 0.046 0.052 0.042 0.048 0.046 0
100 0.052 0.038 0.046 0.058 0.042 0.054 0.048 0.062 0.046 0
110 0.040 0.034 0.038 0.032 0.040 0.052 0.046 0.042 0.034 0
120 0.056 0.044 0.066 0.046 0.046 0.066 0.058 0.052 0.038 0
130 0.052 0.058 0.054 0.056 0.050 0.058 0.052 0.068 0.050 0
140 0.050 0.060 0.054 0.042 0.040 0.074 0.068 0.062 0.058 0
150 0.056 0.064 0.054 0.058 0.060 0.054 0.052 0.042 0.046 0
Table 10: Empirical rejection rates under $H_{1}$ in Experiment IV
Smooth tests SW JB KS
$m$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\widehat{K}\&\chi^{2}_{1}$ $\widehat{K}\&H(x)$
10 0.040 0.998 0.998 0.996 0.994 0.998 0.998 0.994 0.896 0.036
20 0.034 1 1 1 1 1 1 1 1 0.394
30 0.034 1 1 1 1 1 1 1 1 0.834
40 0.024 1 1 1 1 1 1 1 1 0.978
50 0.036 1 1 1 1 1 1 1 1 1
60 0.016 1 1 1 1 1 1 1 1 1
70 0.050 1 1 1 1 1 1 1 1 1
80 0.042 1 1 1 1 1 1 1 1 1
90 0.032 1 1 1 1 1 1 1 1 1
100 0.018 1 1 1 1 1 1 1 1 1
110 0.032 1 1 1 1 1 1 1 1 1
120 0.040 1 1 1 1 1 1 1 1 1
130 0.028 1 1 1 1 1 1 1 1 1
140 0.028 1 1 1 1 1 1 1 1 1
150 0.028 1 1 1 1 1 1 1 1 1
Table 11: Empirical frequency of $\widehat{K}$ in Experiment IV
Under $H_{0}$ Under $H_{1}$
$m$ $\widehat{K}=1$ $\widehat{K}=2$ $\widehat{K}=3$ $\widehat{K}=4$ $\widehat{K}=5$ $\widehat{K}=1$ $\widehat{K}=2$ $\widehat{K}=3$ $\widehat{K}=4$ $\widehat{K}=5$
10 0.972 0.024 0.004 0 0 0 0.990 0.008 0.002 0
20 0.986 0.012 0 0.002 0 0 0.982 0.018 0 0
30 0.980 0.018 0.002 0 0 0 0.990 0.008 0 0.002
40 0.992 0.008 0 0 0 0 0.994 0.006 0 0
50 0.990 0.010 0 0 0 0 0.996 0.004 0 0
60 0.996 0.004 0 0 0 0 0.996 0.004 0 0
70 0.994 0.006 0 0 0 0 0.994 0.006 0 0
80 0.996 0.004 0 0 0 0 0.996 0.004 0 0
90 0.996 0.004 0 0 0 0 1 0 0 0
100 0.990 0.010 0 0 0 0 0.998 0.002 0 0
110 0.994 0.006 0 0 0 0 0.996 0.004 0 0
120 0.992 0.006 0.002 0 0 0 1 0 0 0
130 0.992 0.008 0 0 0 0 1 0 0 0
140 0.986 0.014 0 0 0 0 0.988 0.012 0 0
150 0.998 0.002 0 0 0 0 1 0 0 0
6 An empirical application
In this section, we illustrate the proposed smooth tests using data from the OECD Programme for International Student Assessment (PISA) 2018. PISA is a large-scale international assessment coordinated by the Organisation for Economic Co-operation and Development (OECD) that evaluates educational systems by measuring students’ learning outcomes and school characteristics across countries (schleicher2019pisa). The dataset used in this study is publicly available at https://www.kaggle.com/datasets/dilaraahan/pisa-2018-school-questionnaire. Since PISA data are widely used to compare educational resources across countries, statistical inference for group means is commonly employed, the validity of which depends crucially on distributional assumptions such as normality. However, educational indicators, including school size and student–teacher ratios, often exhibit substantial cross-country heterogeneity and dispersion, and prior studies have documented skewed distributions in school-level resource measures (hanushek2011economics; chandir2022student). Deviations from normality may therefore affect variance estimation and the reliability of ANOVA-based inference, making it important to formally assess this assumption.
We consider two response variables, STRATIO (student–teacher ratio) and SCHSIZE (school size), measured at the school level, and use CNTRYID (country identifier) to define the grouping structure. Table 12 summarizes key sample characteristics after removing missing values, including the total number of observations $N$, the number of groups $J$, and the minimum group size $\min_{j}N_{j}$. The resulting group structure satisfies the conditions required by the asymptotic framework in Section 3 (see Table 1), thereby supporting the application of the proposed models and normality tests. For each response variable (STRATIO and SCHSIZE) under the grouping structure defined by CNTRYID, we fit the ANOVA models (3.1), (3.9), and (3.10), and apply the corresponding smooth tests for normality. For each model, we consider two implementations of the smooth test: one with fixed $K=4$, and one based on the data-driven procedure with $D=5$.
Table 12: Summary of the response and grouping variables and sample characteristics in the PISA 2018 dataset
Response variable Grouping variable NN $J$ $\min_{j}N_{j}$
STRATIO CNTRYID 18042 76 39
SCHSIZE CNTRYID 18321 76 42
The results show that, for both response variables, STRATIO and SCHSIZE, the null hypothesis of normality is rejected at the $0.1\%$ significance level across all three model specifications and both implementations of the proposed tests. This uniform rejection pattern indicates strong robustness of the findings. Overall, there is compelling evidence that the distributions of STRATIO and SCHSIZE deviate substantially from normality. These deviations have important implications for PISA data, where ANOVA models are commonly used to compare educational resources across countries. For variables such as student–teacher ratios and school size, pronounced skewness or heavy-tailed behavior likely reflects substantial heterogeneity in educational systems, with a small number of countries or schools exhibiting extreme values. In such settings, classical $F$-tests may be sensitive to these distributional features, potentially leading to distorted inference when comparing group means. Consequently, ANOVA-based comparisons of educational resources should be interpreted with caution, and more robust approaches, such as transformations, rank-based methods, or other distribution-free procedures, may be preferable in practice. More broadly, these findings underscore the importance of formally assessing distributional assumptions prior to conducting inference in cross-country educational studies.
@@ 7 Concluding remarks
In this paper, we propose and examine data-driven Neyman’s smooth tests for assessing the normality assumption in ANOVA models with a potentially diverging number of groups. For three types of one-way fixed effects models, we derive the asymptotic properties of the proposed tests and validate their finite-sample performance through extensive numerical studies. Our results provide a rigorous and practical tool for evaluating normality in a broad range of ANOVA settings.
Several directions remain open for future research. First, while the proposed tests are developed for specific ANOVA models, the true data-generating mechanism may not be known in practice. A natural extension is to combine our procedures with preliminary structural tests to identify the most appropriate ANOVA specification before applying the proposed methodology. Second, the assumption of independent random errors can be relaxed. For example, repeated clinical trials often involve correlated errors (ma2012beyond; schober2018repeated; langenberg2022repeated). In such cases, our smooth test framework may be adapted by incorporating estimated correlation structures. Finally, for the selection rule (4.1), it is of theoretical interest to explore the scenario where the upper bound $D=D(N)$ diverges slowly with $N$, that is, $D\to\infty$ as $N\to\infty$.
A.1 Proof of Theorem 1
We first study the estimation effects of $\widehat{\mu}-\mu$ and $\widehat{\sigma}^{2}-\sigma^{2}$. Note that
$\widehat{\mu}-\mu=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\mu\right)=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\varepsilon_{ij}=\frac{\sigma}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}.$ (A.1)
Under $H_{0}$, $e_{ij}\sim\mathcal{N}(0,1)i.i.d$., then $\sqrt{N}(\widehat{\mu}-\mu)\sim\mathcal{N}(0,\sigma^{2})$, and $\widehat{\mu}-\mu=O_{p}(N^{-1/2})$ as $N\to\infty$. Besides,
$\widehat{\sigma}^{2}-\sigma^{2}=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\left(Y_{ij}-\widehat{\mu}\right)^{2}-\sigma^{2}\right)$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\varepsilon_{ij}^{2}-\sigma^{2}\right)-\left(\widehat{\mu}-\mu\right)^{2}$